y-3=-2(x+1)
what is the slope of the line represented?

what ordered pair is this point-slope equation based on?
what is the equation in slope-intercept form?

Answer by JBarnum(2146)

this is not a percentage problem
y-3=-2(x+1)
y=-2x-2+3
y=-2x+1
y=mx+b
m=slope=-2
b=intercept=1